### v1

@@ Abstract
Methods. We first characterise intrinsic time delays in BL Lacs and Flat Spectrum Radio Quasars in standard one-zone time-dependent synchrotron self-Compton or external Compton models, during flares produced by particle acceleration and cooling processes. We simulate families of flares with both intrinsic and external LIV-induced energy-dependent delays. Discrimination between intrinsic and LIV delays is then investigated in two different ways. A technique based on Euclidean distance calculation between delays obtained in the synchrotron and in the inverse-Compton spectral bumps is used to assess their degree of correlation. A complementary study is performed using spectral hardness versus intensity diagrams in both energy ranges.
Results. We show that the presence of non-negligible LIV effects, which essentially act only at very high energies (VHE), can drastically reduce the strong correlation expected between the X-ray and the VHE gamma-ray emission in leptonic scenarios. The LIV phenomenon can then be hinted at measuring the Euclidean distance $d_{E}$ from simultaneous X-ray and gamma-ray flare monitoring. Large values of minimal distance $d_{E,min}$ would directly indicate the influence of non-intrinsic time delays possibly due to LIV in SSC flares. LIV effects can also significantly modify the VHE hysteresis patterns in hardness-intensity diagrams and even change their direction of rotation as compared to the X-ray behaviour. Both observables could be used to discriminate between LIV and intrinsic delays, provided high quality flare observations are available.
@@ 1 Introduction
One of the most challenging ambition of modern physics is to unify the realms described by quantum field theory and general relativity. This unification could be achieved through a still in the making unique theory of quantum gravity (QG). The main problem encountered in the process of developing this new fundamental theory is the lack of experimental input, due to the fact the energy scale involved in quantum gravity theories is believed to be extremely high, of the order of the Planck energy, $E_{P}\sim 10^{19}$ GeV. Moreover, extracting observable predictions from yet tentative models has been notoriously difficult. For a more detailed discussion about this topic as well as a detailed review of the state of the art of the different theoretical and experimental efforts currently on-going in QG phenomenology, the reader is invited to refer to Addazi et al. (2022).
Nevertheless, a Lorentz invariance violation (LIV) or deformation appears as a possible effect that could be tested with present-day gamma-ray experiments. These effects are usually represented by a modified dispersion relation for photons in vacuum of the form:
@@
where $E$ and $p$ are the photon energy and momentum, $c$ is the standard (low energy) speed of light, $n$ is the correction order, and $E_{QG,n}$ the so-called quantum gravity energy scale, which may be different for different values of $n$. The correction terms are expected to become significant for $E\sim E_{QG,n}\sim E_{P}$, but should be present at lower energies. The $\pm$ sign accounts for both subluminal and superluminal effects. The modified dispersion relation (1) leads to energy-dependent velocities inducing delays $\Delta t$ in the arrival time of gamma-ray photons with different energies, assuming they are emitted at the same time and location:
$\Delta t=\pm\frac{n+1}{2}\frac{\Delta E^{n}}{E_{QG,n}^{n}}\times\kappa_{n}(z),$ (2)
with $\Delta E^{n}=|E^{n}_{1}-E^{n}_{2}|$ the energy difference at order $n$, $z$ the redshift and $\kappa_{n}(z)$ an increasing function of redshift accounting for cosmological effects. Variation with cosmic distance then appears as a distinctive feature of QG-induced delays. Instead of working with $\Delta t$, a common practice is to use the time delay over energy difference, noted $\tau_{n}$, and defined as:
@@
where $H_{0}$ is the Hubble constant. In the following, we will deal only with this quantity, but for completeness, let us add that $\kappa_{n}$ has been mostly taken from Jacob & Piran (2008) so far, where the expression
@@
is obtained, with $\Omega_{m}$ and $\Omega_{\Lambda}$ respectively the matter density and the dark energy density of the Universe. In the context of gamma-ray astronomy, the quantity $\tau_{n}$ is usually given in units of s TeV-n or s GeV-n. Provided very high energy photons propagate through very long distances, such delays could be detectable with present-day experiments, at least for a linear correction to the dispersion relation ($n=1$). We will only consider this case in the following, noting $\tau\equiv\tau_{n=1}$, expressed in s TeV-1.
@@
It has to be noted that the above description sticks to simple cases when considering LIV delays. In particular, Eq. 4 is only one possible way to account for distance in LIV delay computation (see e.g. Rosati et al., 2015) and other possible QG-related LIV effects, such as the modification of Extra-galactic Background Light (EBL) absorption (see Addazi et al., 2022, §5.3), have been neglected. Moreover, classical intrinsic effects arising in the source can limit the possibility of having bunches of gamma-ray photons emitted from the same location and at the same time which is the primary focus of the present paper.
@@
Following Eq. 3, a strategy currently in use to search for LIV signatures is to look for energy-dependent time delays in the gamma-ray signal coming from remote and variable cosmic sources such as flaring active galactic nuclei (AGN), pulsars and gamma-ray bursts (GRB, Addazi et al., 2022; Terzić et al., 2021). These sources are detected at MeV–GeV energies by satellites and at TeV energies with ground-based gamma-ray instruments such as imaging atmospheric Cherenkov telescopes (IACTs). However, measured delays can include time lags that can be generated by classical radiative processes at the sources, or any other effect that could modify photon propagation depending on its energy. The latter effects can be due, for example (but not only), to exotic plasma effects at the source (Latorre et al., 1995; Myers & Pospelov, 2003; Albert et al., 2008), or to the delayed production of secondary photons in cascades deflected by extragalactic magnetic fields (Taylor et al., 2011). In the present paper, we focus on the former, the so-called “source intrinsic effects”, specifically dealing with the case of blazar flares. Such classical intrinsic effects can lead to energy-dependent time delays, which have been shown to be non-negligible at very high energies (VHE) in standard flare models (Perennes et al., 2020) for a large domain of initial parameters, and would need to be discriminated from LIV-induced delays in the case of a significant lag detection.
@@
Up to now, no significant lag was measured and limits on $E_{QG}$ were derived, except in one occasion for a flare of the blazar Mkn~501 recorded by MAGIC in July 2005 (Albert et al., 2007). It is however probable that new generation of instruments with improved sensitivity will lead to more frequent lag detection. At this point, it will be necessary to disentangle between source intrinsic effects and LIV-induced delays.
@@
A first way to achieve that goal relies on the fact source intrinsic time delays are not expected to directly depend on the distance of the source. On the contrary, as described by Eq. 3, the LIV time delay has an explicit dependence on the redshift, expressed by the function $\kappa_{n}(z)$ which is continuously increasing with $z$. So it is expected that detected delays should significantly vary with $z$ if they are due to LIV, while they have no immediate reasons to do so if they are due to intrinsic effects. Indeed, it is the variability time scale of the flares (typically from minutes to hours) which allow intrinsic delay detection, and determines (by the causality argument) the physical size of the emitting zone. This size directly governs, in turn, the particle acceleration and cooling timescales and intrinsic time delays. So in a typical sample of blazar flares, the observed intrinsic delays are not supposed to depend primarily on distance  [^1] At second order, the possibility of a complex Malmquist bias, taking into account Doppler boosting effects, would deserve a further analysis, beyond the scope of the present study.. Therefore, a discrimination strategy relies on population studies involving a large number of sources over a broad range of redshifts. Such analyses have already been published using multiple GRBs observed by satellites (see e.g. Ellis et al., 2019, and references therein), leading to lower limits on $E_{QG}$ up to $10^{17}$ GeV for $n=1$, and an effort is currently on-going to perform a population study with all sources (AGNs, GRBs, and pulsars) detected by present-day IACT arrays (Bolmont et al., 2022).
@@
Another, yet complementary way to separate intrinsic and LIV-induced delays relies on the study of the intrinsic effects through the modelling of source emission mechanisms. A first attempt in that direction was presented by Perennes et al. (2020) for flares of BL Lac sources. Starting from a simple one-zone purely leptonic Synchrotron Self-Compton (SSC) scenario, and using conservative physical parameters to describe the emitting plasma, the authors showed how energy-dependent intrinsic lags are produced during single flares generated by acceleration and cooling processes. They identified two main regimes, depending on the domain of parameters, one where the lags increase with energy at VHE (with lower energies leading higher ones), and one where the lags decrease with energy at VHE (with higher energies leading lower ones). In the first ”acceleration-driven” regime, flares are globally dominated by a relatively slow acceleration mechanism and reach their peak before electrons reach their maximal energy, while in the second ”cooling-driven” regime they are globally dominated by the cooling effect and reach their peak after electrons reach their maximal energy. These intrinsic lags can be large at VHE, reaching several hundreds seconds per TeV, meaning that present-day observations could already been used to constrain this kind of models, at least for very bright flares. It is also found that intrinsic lags vary with time during the flare. Since LIV delays obviously do not, this may offer another way to separate LIV and intrinsic delays. However, measuring the evolution of the delay in VHE during a single BL Lac flare seems completely out of reach at the moment, and may even be challenging with the next generation of instruments such as the Cherenkov Telescope Array  [^2] https://www.cta-observatory.org (CTA).
@@
In this paper we explore new ways to possibly disentangle potential LIV lags from intrinsic delays. The results shown were obtained with a software library called AGNES  [^3] The code can be obtained by request to the authors. (AGN Evolution Simulator) specifically developed for that purpose by C. Levy. In Sect. 2, we recall and further develop the study by Perennes et al. (2020), examining the impact on intrinsic delays of adiabatic expansion and external inverse-Compton emission phenomena expected in Flat Spectrum Radio Quasars (FSRQs). In Sect. 3, LIV effects are added to the emission model output, allowing the simultaneous description and direct comparison of the two types of delays. We show how multi-wavelength monitoring of blazar flares, covering X-ray and gamma-ray domains, could offer direct possibilities to distinguish the presence of an additional lag potentially due to LIV over classical intrinsic delays even in the case where only one source is considered at one particular redshift. Two techniques are proposed to combine simultaneous X-ray and gamma-ray data with the goal to discriminate LIV from intrinsic effects. These techniques are tested on the model output. Conclusions and perspectives are given in Sect. 4.
@@
2 Characterisation of intrinsic time delays at VHE in blazar
@@
Several time dependent non-thermal emission models for radio, X-ray and gamma-ray flares of AGN proposed in the literature lead to the existence of intrinsic time delays generated from radiation mechanisms, for instance in Blandford & Königl (1979); Marscher & Gear (1985); Celotti et al. (1991); Kirk et al. (1998); Katarzyński et al. (2003); Lewis et al. (2016). Such types of leptonic models are extensively used in the literature to describe the temporal evolution of flares generated by particle injection, acceleration and cooling. They are fairly constrained and are considered as reference models by the community since they can reproduce simple blazar flares reasonably well, at least for currently available datasets. For VHE flares varying quickly on timescale $t_{0}$, the size of the emission zone is limited to $R\leq ct_{0}\delta_{b}/(1+z)$ by the causality argument, where $\delta_{b}$ is the Doppler boost of the blob, and emission models dominated by one compact zone are often preferred. We therefore adopt here a leptonic scenario to further analyse the intrinsic time delays at high and very high energies, describing the VHE emission zone as a single optically thin small plasma blob travelling along the blazar’s jet with a bulk Lorentz factor $\Gamma$, and harbouring an electron-positron population submitted to acceleration and cooling processes.
@@ 2.1 Leptonic flare scenario
Here $t$ and $\gamma$ are the time and the individual particle Lorentz factor. $C_{\mathrm{acc}}(t)\propto A(t)$ is the acceleration term, with $A(t)$ a function defined as:
@@
with $A_{0}$ the initial acceleration amplitude and $m_{a}$ the evolution index. The radiative term $C_{\mathrm{rad}}(t)$ is defined as $C_{\mathrm{rad}}(t)\equiv C^{\mathrm{SSC}}_{\mathrm{rad}}(t)+C^{ext}_{\mathrm{rad}}$, where $C^{ext}_{\mathrm{rad}}$ stands for the particle radiation and cooling due to inverse-Compton losses on any external photon field, here assumed to be constant. The other term $C^{\mathrm{SSC}}_{\mathrm{rad}}(t)\propto B(t)^{2}$ accounts for the particle radiation and cooling due to the synchrotron losses and the synchrotron-self-Compton losses, here assumed smaller than and proportional to synchrotron losses. The evolution of the magnetic field $B(t)$ is parameterised as
@@
with $B_{0}$ the initial magnetic field strength, and $m_{b}$ the evolution index. $t_{0}$ is the characteristic evolution time, taken as the time needed for a sound wave to propagate through the blob of relativistic plasma of radius $R$. $C_{\mathrm{adia}}(t)$ in Eq. 5 describes the effect of slow adiabatic expansion (or compression) of the compact spherical emission zone with radius $R(t)$ parameterised as
@@
with $R_{0}$ the initial blob’s radius, and $m_{r}$ the evolution index. $N^{*}(t,\gamma)$ is the number density per unit Lorentz factor $\gamma$, modified to take into account the blob’s radius evolution index, which reads
@@
Figure 1: Distributions obtained from solving the differential equation governing the lepton number density, for the set of reference input parameters given in Table 1. No stationary state contribution is included here. Left: lepton spectra temporal evolution. Centre: spectral energy distributions temporal evolution generated as seen by the observer. Right: normalised multi-wavelength light curves.
In BL Lacs, SSC radiation is expected to dominate in the absence of bright accretion disk, broad line region (BLR) and dust torus. Introducing adiabatic processes and external Inverse-Compton (EIC) emission allows for further exploration beyond the BL Lac case analysed by Perennes et al. (2020), especially extending it to FSRQs, where the EIC emission should play a significant role. The additional number of photons brought by the external radiation fields contribute to generate VHE photons and boost the flux in the SED inverse Compton component. External photon fields generated by bright disk, BLR and torus in FSRQ are expected to dominate the EIC component respectively one after the other with the increase in the distance $d$ of the emitting blob relative to the central black hole, and can be roughly modelled with a black-body emission centred on a specific temperature. To illustrate their effect on intrinsic time delays, we consider here a static photon field, generated by the accretion disc and reprocessed by the clouds of the BLR located at the distance $d$ with typically $0.01\ \mathrm{pc}<d<0.1\ \mathrm{pc}$. This results in an ambient radiation field which is Doppler boosted as seen in the blob’s frame when moving towards the BLR clouds. The cooling coefficient $C^{\mathrm{ext}}_{\mathrm{rad}}$ is taken to be proportional to the energy density $U_{\mathrm{ext}}$ as described in Inoue & Takahara (1996):
@@
with $\Gamma_{b}$ the Lorentz factor of the plasma blob, $L_{AD}$ the accretion disc luminosity, $d$ the distance between the accretion disc and the plasma blob, and $\tau$ the reprocessing efficiency combined with the covering factor of the BLR. A significant fraction of the external photons (those that are directed or reflected head-on towards the relativistic plasma blob) are strongly blue-shifted in the source’s rest frame. In particular, photons reprocessed by the BLR can reach keV energies (X-rays) and therefore be boosted up to TeV energies through inverse Compton interactions with the blob’s leptons. For illustrative purposes, we adopt hereafter a typical value for the reprocessed luminosity $\tau L_{AD}=10^{41}$ erg s-1 and only vary the distance $d$ from 0.03 pc to infinity, to analyse stronger and weaker EIC effects. Note that in reality at large distances (typically $d>0.1$ pc), once the blob has crossed the BLR, the ambient radiation due to the dust torus is supposed to dominate the EIC component. Here we explore only the qualitative effect of EIC emissions on the temporal time delays, without aiming to reproduce in detail a specific AGN.

### v2

@@ Abstract
Methods. We first characterised the intrinsic time delays in BL Lacs and flat-spectrum radio quasars in standard one-zone time-dependent synchrotron self-Compton or external Compton models, during flares produced by particle acceleration and cooling processes. We simulated families of flares with both intrinsic and external LIV-induced energy-dependent delays. Discrimination between intrinsic and LIV delays is then investigated in two different ways. A technique based on Euclidean distance calculation between delays obtained in the synchrotron and in the inverse-Compton spectral bumps is used to assess their degree of correlation. A complementary study is performed using spectral hardness versus intensity diagrams in both energy ranges.
Results. We show that the presence of non-negligible LIV effects, which essentially act only at very high energy (VHE), can drastically reduce the strong correlation expected between the X-ray and the VHE gamma-ray emission in leptonic scenarios. The LIV phenomenon can then be hinted at measuring the Euclidean distance $d_{E}$ from simultaneous X-ray and gamma-ray flare monitoring. Large values of minimum distance $d_{E,min}$ would directly indicate the influence of non-intrinsic time delays possibly due to LIV in SSC flares. LIV effects can also significantly modify the VHE hysteresis patterns in hardness-intensity diagrams and even change their direction of rotation compared to the X-ray behaviour. Both observables could be used to discriminate between LIV and intrinsic delays, provided high-quality flare observations are available.
@@ 1 Introduction
One of the most challenging ambitions of modern physics is to unify the realms described by quantum field theory and general relativity. This unification could be achieved through a still-in-the-making unique theory of quantum gravity (QG). The main problem encountered in the process of developing this new fundamental theory is the lack of experimental input, due to the fact that the energy scale involved in quantum gravity theories is believed to be extremely high, of the order of the Planck energy, $E_{P}\sim 10^{19}$ GeV. Moreover, extracting observable predictions from still tentative models has been notoriously difficult. For a more detailed discussion about this topic, as well as a detailed review of the state of the art of the different theoretical and experimental efforts currently ongoing in QG phenomenology, we refer to Addazi et al. (2022).
Nevertheless, a Lorentz invariance violation (LIV) or deformation appears as a possible effect that could be tested with present-day gamma-ray experiments. These effects are usually represented by a modified dispersion relation for photons in vacuum of the form
@@
where $E$ and $p$ are the photon energy and momentum, $c$ is the standard (low-energy) speed of light, $n$ is the correction order, and $E_{QG,n}$ is the quantum gravity energy scale, which may be different for different values of $n$. The correction terms are expected to become significant for $E\sim E_{QG,n}\sim E_{P}$, but should be present at lower energies. The $\pm$ sign accounts for both subluminal and superluminal effects. The modified dispersion relation (1) leads to energy-dependent velocities inducing delays $\Delta t$ in the arrival time of gamma-ray photons with different energies, assuming they are emitted at the same time and location:
$\Delta t=\pm\frac{n+1}{2}\frac{\Delta E^{n}}{E_{QG,n}^{n}}\times\kappa_{n}(z).$ (2)
Here $\Delta E^{n}=|E^{n}_{1}-E^{n}_{2}|$ is the energy difference at order $n$, $z$ the redshift, and $\kappa_{n}(z)$ an increasing function of redshift accounting for cosmological effects. Variation with cosmic distance then appears as a distinctive feature of QG-induced delays. Instead of working with $\Delta t$, a common practice is to use the time delay over energy difference, denoted $\tau_{n}$, and defined as
@@
where $H_{0}$ is the Hubble constant. In the following we deal only with this quantity, but for completeness, we note that $\kappa_{n}$ has been mostly taken from Jacob & Piran (2008) so far, where the expression
@@
is obtained, with $\Omega_{m}$ and $\Omega_{\Lambda}$ respectively the matter density and the dark energy density of the Universe. In the context of gamma-ray astronomy, the quantity $\tau_{n}$ is usually given in units of s TeV-n or s GeV-n. Provided very high-energy photons propagate through very long distances, such delays could be detectable with present-day experiments, at least for a linear correction to the dispersion relation ($n=1$). We only consider this case in the following, noting $\tau\equiv\tau_{n=1}$, expressed in s TeV-1.
@@
It should be noted that the above description sticks to simple cases when considering LIV delays. In particular, Eq. 4 is only one possible way to account for distance in LIV delay computation (see e.g. Rosati et al., 2015; Pfeifer, 2018) and other possible QG-related LIV effects, such as the modification of extra-galactic background light (EBL) absorption (see Addazi et al., 2022, §5.3), are neglected. Moreover, classical intrinsic effects arising in the source can limit the possibility of having bunches of gamma-ray photons emitted from the same location and at the same time, which is the primary focus of the present paper.
@@
Following Eq. 3, a strategy currently in use to search for LIV signatures is to look for energy-dependent time delays in the gamma-ray signal coming from remote and variable cosmic sources, such as flaring active galactic nuclei (AGN), pulsars, and gamma-ray bursts (GRBs, Addazi et al., 2022; Terzić et al., 2021). These sources are detected at MeV–GeV energies by satellites and at TeV energies with ground-based gamma-ray instruments, such as imaging atmospheric Cherenkov telescopes (IACTs). However, measured delays can include time lags that can be generated by classical radiative processes at the sources, or any other effect that could modify photon propagation depending on its energy. The latter effects can be due, for example (but not only), to exotic plasma effects at the source (Latorre et al., 1995; Myers & Pospelov, 2003; Albert et al., 2008) or to the delayed production of secondary photons in cascades deflected by extragalactic magnetic fields (Taylor et al., 2011). In the present paper, we focus on the former, called the source intrinsic effects, specifically dealing with the case of blazar flares. Such classical intrinsic effects can lead to energy-dependent time delays, which have been shown to be non-negligible at very high energy (VHE) in standard flare models (Perennes et al., 2020) for a large domain of initial parameters, and would need to be discriminated from LIV-induced delays in the case of a significant lag detection.
@@
To date, no significant lag has been measured or limits on $E_{QG}$ derived, except on one occasion for a flare of the blazar Mkn~501 recorded by MAGIC in July 2005 (Albert et al., 2007). It is probable, however, that a new generation of instruments with improved sensitivity will lead to more frequent lag detection. At this point, it will be necessary to disentangle between source intrinsic effects and LIV-induced delays.
@@
The first way to achieve that goal relies on the fact that source intrinsic time delays are not expected to directly depend on the distance of the source. On the contrary, as described by Eq. 3, the LIV time delay has an explicit dependence on the redshift, expressed by the function $\kappa_{n}(z),$ which continuously increases with $z$. So it is expected that detected delays should significantly vary with $z$ if they are due to LIV, while they have no immediate reasons to do so if they are due to intrinsic effects. It is the variability timescale of the flares (typically from minutes to hours) that allows intrinsic delay detection, and determines (by the causality argument) the physical size of the emitting zone. This size directly governs, in turn, the particle acceleration and cooling timescales and intrinsic time delays. So in a typical sample of blazar flares, the observed intrinsic delays are not supposed to depend primarily on distance. [^1] At second order, the possibility of a complex Malmquist bias, taking into account Doppler boosting effects, would deserve a further analysis, beyond the scope of the present study. Therefore, a discrimination strategy relies on population studies involving a large number of sources over a broad range of redshifts. Similar analyses have already been published using multiple GRBs observed by satellites (see e.g. Ellis et al., 2019, and references therein), leading to lower limits on $E_{QG}$ up to $10^{17}$ GeV for $n=1$, and an effort is currently ongoing to perform a population study with all sources (AGN, GRBs, and pulsars) detected by present-day IACT arrays (Bolmont et al., 2022).
@@
Another complementary way to separate intrinsic and LIV-induced delays relies on the study of the intrinsic effects through the modelling of source emission mechanisms. A first attempt in that direction was presented by Perennes et al. (2020) for flares of BL Lac sources. Starting from a simple one-zone purely leptonic synchrotron self-Compton (SSC) scenario, and using conservative physical parameters to describe the emitting plasma, the authors showed how energy-dependent intrinsic lags are produced during single flares generated by acceleration and cooling processes. They identified two main regimes, depending on the domain of parameters, one where the lags increase with energy at VHE (with lower energies leading to higher ones), and one where the lags decrease with energy at VHE (with higher energies leading to lower ones). In the first regime (called acceleration-driven), flares are globally dominated by a relatively slow acceleration mechanism and reach their peak before electrons reach their maximum energy, while in the second regime (called cooling-driven), they are globally dominated by the cooling effect and reach their peak after electrons reach their maximum energy. These intrinsic lags can be large at VHE, reaching several hundreds of seconds per TeV, meaning that present-day observations could already be used to constrain this kind of model, at least for very bright flares. It is also found that intrinsic lags vary with time during the flare. Since LIV delays obviously do not, this may offer another way to separate LIV and intrinsic delays. However, measuring the evolution of the delay in VHE during a single BL Lac flare seems completely out of reach at the moment, and may even be challenging with the next generation of instruments such as the Cherenkov Telescope Array  [^2] https://www.ctao.org (CTA).
@@
In this paper we explore new ways to possibly disentangle potential LIV lags from intrinsic delays. The results shown were obtained with a software library called the AGN Evolution Simulator (AGNES),  [^3] The code can be downloaded from https://gitlab.in2p3.fr/julien.bolmont/AGNES. specifically developed for that purpose by C. Levy. In Sect. 2 we recall and further develop the study by Perennes et al. (2020), examining the impact on intrinsic delays of adiabatic expansion and external inverse-Compton emission phenomena expected in flat-spectrum radio quasars (FSRQs). In Sect. 3 LIV effects are added to the emission model output, allowing the simultaneous description and direct comparison of the two types of delays. We show how multi-wavelength monitoring of blazar flares, covering X-ray and gamma-ray domains, could offer direct possibilities to distinguish the presence of an additional lag potentially due to LIV over classical intrinsic delays, even in the case where only one source is considered at one particular redshift. Two techniques are proposed to combine simultaneous X-ray and gamma-ray data with the goal to discriminate LIV from intrinsic effects. These techniques are tested on the model output. Conclusions and perspectives are given in Sect. 4.
@@
2 Characterisation of intrinsic time delays at very high energy in blazars
@@
Several time dependent non-thermal emission models for radio, X-ray, and gamma-ray flares of AGN proposed in the literature lead to the existence of intrinsic time delays generated from radiation mechanisms (e.g. Blandford & Königl, 1979; Marscher & Gear, 1985; Celotti et al., 1991; Kirk et al., 1998; Katarzyński et al., 2003; Lewis et al., 2016). These types of leptonic models are extensively used in the literature to describe the temporal evolution of flares generated by particle injection, acceleration, and cooling. They are fairly constrained and are considered reference models by the community since they can reproduce simple blazar flares reasonably well, at least for currently available datasets. For VHE flares varying quickly on timescale $t_{0}$, the size of the emission zone is limited to $R\leq ct_{0}\delta_{b}/(1+z)$ by the causality argument, where $\delta_{b}$ is the Doppler boost of the blob, and emission models dominated by one compact zone are often preferred. We therefore adopt here a leptonic scenario to further analyse the intrinsic time delays at high and very high energies, describing the VHE emission zone as a single optically thin small plasma blob travelling along the blazar’s jet with a bulk Lorentz factor $\Gamma$, and harbouring an electron-positron population submitted to acceleration and cooling processes.
@@ 2.1 Leptonic flare scenario
Here $t$ and $\gamma$ are the time and the individual particle Lorentz factor; $C_{\mathrm{acc}}(t)\propto A(t)$ is the acceleration term, with $A(t)$ a function defined as
@@
with $A_{0}$ the initial acceleration amplitude and $m_{a}$ the evolution index. The radiative term $C_{\mathrm{rad}}(t)$ is defined as $C_{\mathrm{rad}}(t)\equiv C^{\mathrm{SSC}}_{\mathrm{rad}}(t)+C^{ext}_{\mathrm{rad}}$, where $C^{ext}_{\mathrm{rad}}$ stands for the particle radiation and cooling due to inverse-Compton losses on any external photon field, here assumed to be constant. The other term $C^{\mathrm{SSC}}_{\mathrm{rad}}(t)\propto B(t)^{2}$ accounts for the particle radiation and cooling due to the synchrotron losses and the synchrotron-self-Compton losses, here assumed smaller than and proportional to synchrotron losses (Katarzyński et al., 2001). The evolution of the magnetic field $B(t)$ is parametrised as
@@
with $B_{0}$ the initial magnetic field strength and $m_{b}$ the evolution index. The parameter $t_{0}$ is the characteristic evolution time, taken as the time needed for a sound wave to propagate through the blob of relativistic plasma of radius $R$. The quantity $C_{\mathrm{adia}}(t)$ in Eq. 5 describes the effect of slow adiabatic expansion (or compression) of the compact spherical emission zone with radius $R(t)$ parametrised as
@@
with $R_{0}$ the blob’s initial radius, and $m_{r}$ the evolution index. The quantity $N^{*}(t,\gamma)$ is the number density per unit Lorentz factor $\gamma$, modified to take into account the blob’s radius evolution index, which reads
@@
Figure 1: Distributions obtained from solving the differential equation governing the lepton number density, for the set of reference input parameters given in Table 1. No stationary state contribution is included here. Left: Lepton spectra temporal evolution. Centre: Spectral energy distributions of temporal evolution, generated as seen by the observer. Right: Normalised multi-wavelength light curves.
In BL Lacs, SSC radiation is expected to dominate in the absence of a bright accretion disc, a broad line region (BLR), and a dust torus. Introducing adiabatic processes and external inverse-Compton (EIC) emission allows further exploration beyond the BL Lac case analysed by Perennes et al. (2020), especially extending it to FSRQs, where the EIC emission should play a significant role. The additional number of photons brought by the external radiation fields contribute to generate VHE photons and boost the flux in the SED inverse-Compton component. External photon fields generated by a bright disc, a BLR, and a torus in FSRQ are expected to dominate the EIC component respectively one after the other with the increase in the distance $d$ of the emitting blob relative to the central black hole, and can be roughly modelled with a blackbody emission centred on a specific temperature. To illustrate their effect on intrinsic time delays, we consider here a static photon field, generated by the accretion disc and reprocessed by the clouds of the BLR located at the distance $d$ with typically $0.01\ \mathrm{pc}<d<0.1\ \mathrm{pc}$. This results in an ambient radiation field that is Doppler boosted as seen in the blob’s frame when moving towards the BLR clouds. The cooling coefficient $C^{\mathrm{ext}}_{\mathrm{rad}}$ is taken to be proportional to the energy density $U_{\mathrm{ext}}$, as described in Inoue & Takahara (1996)
@@
with $\Gamma_{b}$ the Lorentz factor of the plasma blob, $L_{AD}$ the accretion disc luminosity, $d$ the distance between the accretion disc and the plasma blob, and $\tau$ the reprocessing efficiency combined with the covering factor of the BLR. A significant fraction of the external photons (those that are directed or reflected head-on towards the relativistic plasma blob) are strongly blueshifted in the source’s rest frame. In particular, photons reprocessed by the BLR can reach keV energies (X-rays) and therefore be boosted up to TeV energies through inverse-Compton interactions with the blob’s leptons. For illustrative purposes, we adopt hereafter a typical value for the reprocessed luminosity of $\tau L_{AD}=10^{41}$ erg s-1, and only vary the distance $d$ from 0.03 pc to infinity, to analyse stronger and weaker EIC effects. We note that in reality at large distances (typically $d>0.1$ pc), once the blob has crossed the BLR, the ambient radiation due to the dust torus is supposed to dominate the EIC component. Here we explore only the qualitative effect of EIC emissions on the temporal time delays, without aiming to reproduce in detail a specific AGN.